OpenStudy (anonymous):

Marcy has a total of 100 dimes and quarters. If the total value of the coins is $14.05, how many quarters does she have?

OpenStudy (australopithecus):

100 dimes and quarters so dimes = x quarters = y x + y = 100 we note that a dime is 0.10 cents and a quarter is 0.25 cents so x(0.10) + y(0.25) = 14.05 Thus we have two equations x + y = 100 x(0.10) + y(0.25) = 14.05 Do you know how to solve for x and y?

OpenStudy (anonymous):

umm substitution?

OpenStudy (tommo_lcfc):

That is easiest.